OpenStudy (anonymous):

Simplify quantity 4 times x minus 20 all over 4 times x.

OpenStudy (anonymous):

is it equal to zero?

OpenStudy (anonymous):

no i tried that

OpenStudy (anonymous):

(4x-20)/4x =[4(x-5)]/4x =(x-5)/x